OpenStudy (anonymous):

Erica bought 3 1/2 yards of fabric. If she uses 2/3 of the fabric to make a curtain, how much will she have left?

OpenStudy (anonymous):

3 1/2 - (2/3)(3 1/2) = 7/2 - (2/3)(7/2) = 7/2 - 7/3 = 21/6 - 14/6 = 7/6 or 1 1/6 yds

OpenStudy (anonymous):

He will have used 2 1/3 yards. Covert 3 1/2 to an improper fraction and then multiply by 2/3. This would be 7/2 X 2/3 = 14/6 which when converted back to a proper fraction equals 2 1/3. Now, you can subtract the 2 1/3 from the original 3 1/2 yards and you find he still has 1 1/6 yard of fabric

OpenStudy (anonymous):

what are the choices

OpenStudy (anonymous):

you right 1 1/6 is correct thx

OpenStudy (anonymous):

u understand

OpenStudy (anonymous):

yea first i subtracted i did not multiply